OpenStudy (anonymous):

Factor the following using the GCF: 10xy3 - 25x2y2 + 15xy

OpenStudy (anonymous):

im just confused with gcf

OpenStudy (anonymous):

GCF = 5xy

OpenStudy (anonymous):

5xy my man.

OpenStudy (anonymous):

GCF is greatest common factor for numbers you can find it by first factoring them into primes: 10 = 2*5 25 = 5*5 15 = 3*5 5 is common to all 3 so GCF is 5 as for the letters it is the one with lowest power: x , x^2, x - GCF is x y^3, y^2 , y - GCF is y

OpenStudy (mertsj):

Think of it this way: the biggest number that will divide into all the numbers. If it's powers think the biggest power they have in common.